# Active Learning with Task–Driven Representations for Messy Pools

### v1

@@ A.1.3 Models
For F+MNIST, we used the encoder and decoder from Burgess et al., (2017); for CheXpert we used the encoder and decoder from Higgins et al., (2017); for CIFAR-10+100 we used the ResNetVAE used for the CIFAR-10 dataset in Kingma et al., (2016).
@@
For F+MNIST and CheXpert we used the ResNet18 architecture (He et al.,, 2016); for CIFAR-10+100 we used the ResNet50 architecture (He et al.,, 2016).
@@ A.1.4 Active learning
Initial labeled set:
To create our initial labeled set, we randomly sampled 2 labels from “0”, “1”, and “2” for F+MNIST; for CheXpert we randomly sampled 4 labels from “0” and “1”; for CIFAR-10+100 we randomly sampled 2 classes for each of our target classes and “0” class.
@@ A.1.5 Acquisition strategies
For our acquisition strategies, we used EPIG, BALD, Confidence Sampling and the acquisition strategies from our baselines methods which we describe below. To estimate BALD and EPIG, we used the same setup as Bickford Smith et al., (2023): we used 100 realisations of $\theta_{h}$ (for random forests this meant using the individual trees; otherwise this meant sampling from the parameter distribution). For EPIG, we used $M$ samples of $x_{*}$ from a finite set of unlabelled inputs representative of the downstream task, where $M=500$ for CheXpert and F+MNIST, and $M=1000$ for CIFAR-10+100.
@@
We used batch acquisitions for all datasets, with a batch size of 10 for F+MNIST and 100 for the others. We used the “power” batch acquisition strategy from Kirsch et al., (2021) with $\beta=4$ for F+MNIST and $\beta=8$ otherwise. We make this choice as this strategy is both highly scalable and has been shown to give performance comparable to more sophisticated batch acquisition strategies.
@@ A.1.6 Baselines
We used baselines which were specifically designed for scenarios with class imbalance and redundancy (GALAXY (Zhang et al.,, 2022), SIMILAR (Kothawade et al.,, 2021)) or which have shown strong performance in such settings (Cluster Margin (Citovsky et al.,, 2021) from the results in Zhang et al., (2022)). We note that Confidence Sampling (Settles,, 2009), though not treated as a baseline, has also shown strong performance in this setting (as per the results in Zhang et al., (2022)).
@@
For SIMILAR, we used the FLQMI relaxation of the submodular mutual information (SMI) due to memory constraints posed by the FLCMI relaxation. For Cluster Margin, we first experimented with their original hyperparameters ($\epsilon$ such that we have at least 10 clusters and $k_{m}=10k_{t}$). We found this to result in poor performance and so instead adopted the hyperparameters used in Zhang et al., (2022), i.e. choosing $\epsilon$ such that we have exactly 50 clusters and $k_{m}=1.25k_{t}$, where $k_{t}$ is our acquisition batch size.
@@
For a fair comparison with our approach, we used a ResNet18 model for all the baselines on F+MNIST and CheXpert, and a ResNet50 for CIFAR-10+100. We added a final fully–connected hidden layer of 128 hidden units on top of the ResNet models and trained them in a fully supervised fashion. For Cluster Margin, we followed Citovsky et al., (2021) and warm–started the models by training them on a validation set (0.5% of our pool size) that was evenly balanced between the target classes and ‘other’ class; [^1] Note that Cluster Margin still performs worse than our approach despite the unrealistic assumption of a validation set. for SIMILAR, we followed Kothawade et al., (2021) and trained the models fully from scratch; for GALAXY, we followed Zhang et al., (2022) and initialised the ResNet backbones with weights pretrained on ImageNet in a fully supervised fashion. We trained the models using a batch size of 200, the Adam optimiser, and a learning rate of 0.001 for Cluster Margin, SIMILAR and 0.0001 for GALAXY.
@@
For the TRANSFER approach, we replaced our SimCLRv2 encoders pretrained on the pool with: ResNet50 and ResNet101 (He et al.,, 2016) pretrained on ImageNet in a fully supervised fashion, and DinoV2–Small and DinoV2–Big (Oquab et al.,, 2023) pretrained in a self–supervised fashion on the curated dataset discussed in in Oquab et al., (2023) [^2] The ResNet models were accessed through TorchVision and the Dino–V2 models were accessed from the PyTorch Hub (Paszke et al.,, 2019). Specifically, DinoV2–Small corresponds to facebookresearch/dinov2/dinov2_vits14 and DinoV2–Big corresponds to facebookresearch/dinov2/dinov2_vitb14..
@@
Dino–V2 models:
Similar to Oquab et al., (2023), we followed a similar finetuning setup to the one in Touvron et al., (2022). Specifically, we used a batch size of 512, the Adam optimiser, learning rate of $3\times 10^{-4}$ with 5 epochs warmup and cosine decay, and label smoothing with level 0.1. We used the same augmentations in Section A.1.2.
@@ Appendix B Additional plots
B.1 Full active learing curves for Table 1
@@
Figure 5: Test accuracy for our TD-SPLIT approach on F+MNIST, CIFAR-10+100, CheXpert and the baselines considered in Table 1. All experiments were run for 4 seeds. Solid line shows mean and shading $\pm 1$ standard error.
@@
Figure 6: Test accuracy for our $S$-FT approach on F+MNIST, CIFAR-10+100, CheXpert and the baselines considered in Table 1. All experiments were run for 4 seeds. Solid line shows mean and shading $\pm 1$ standard error.
@@ C.3 Different prediction heads
To show that our approach is compatible with different prediction heads, we replaced the random forest prediction head with a 1 layer neural network with 128 hidden units. We used the Laplace approximation MacKay, 1992a to infer the parameter distribution, where we used a standard Gaussian prior, $\mathcal{N}(0,I)$, and a diagonal, tempered posterior (Aitchison,, 2020), with tempering implemented by raising the likelihood term to a power of $\text{dim}(\theta_{h})$ (i.e. the parameter count of the prediction head).
@@ Appendix D Additional results
In this section, we provide additional results about the computational cost of our approach, using our task–driven representations for the baselines and the acquisition counts for different approaches.
@@ D.1 Task–driven representations improve baselines
We see that integrating the baselines into our approach significantly boosts their performance on all dataset. We note, however, that their performance is still lagging behind our approach with the EPIG acquisition strategy.
@@
SIMILAR (Kothawade et al.,, 2021) $93.82\pm 0.18$ $30.87\pm 1.57$ $71.94\pm 0.47$
GALAXY (Zhang et al.,, 2022) $84.74\pm 0.74$ $55.28\pm 0.42$ $78.76\pm 0.35$
Cluster Margin (Citovsky et al.,, 2021) $94.24\pm 0.17$ $32.79\pm 0.45$ $75.41\pm 0.29$
US+EPIG (SimCLRv2, Bickford Smith et al., (2024)) $98.53\pm 0.12$ $6.19\pm 0.42 77$$7.84\pm 0.28$
US+EPIG (VAE, Bickford Smith et al., (2024)) $94.50\pm 0.34$ $30.47\pm 1.17$ $66.69\pm 0.56$
@@
TD-SIMILAR (Kothawade et al.,, 2021) $98.05\pm 1.19$ $9.90\pm 0.23 77$$7.26\pm 1.06$
TD-GALAXY (Zhang et al.,, 2022) $99.30\pm 0.13$ $73.62\pm 0.77$ $78.07\pm 0.73$
TD-Cluster Margin (Citovsky et al.,, 2021) $99.32\pm 0.18$ $70.47\pm 2.58$ $76.30\pm 0.38$
@@ D.2 Acquisition counts for different approaches
From Table 5, we note two things. First we note that using our approach, whether that is with the baselines or instead of unsupervised representations, improves the count of the target classes by a large margin. This suggests that the gains displayed in Table 4 are not merely from using a better model, but also from making better acquisitions. Secondly, we note that the best performing method (TD-FT) does not have the largest amount of target classes acquired. This suggests that the precise data point acquired is important, not just its clas (Yang et al.,, 2023).
@@
SIMILAR (Kothawade et al.,, 2021) $486\pm 6$ $920\pm 237$
GALAXY (Zhang et al.,, 2022) $118\pm 13$ $1806\pm 32$
Cluster Margin (Citovsky et al.,, 2021) $32\pm 1$ $1287\pm 53$
US+EPIG (SimCLRv2, Bickford Smith et al., (2024)) $119\pm 8$ $936\pm 35$
US+EPIG (VAE, Bickford Smith et al., (2024)) $118\pm 4$ $529\pm 17$
TD-SIMILAR (Kothawade et al.,, 2021) $493\pm 1$ $2231\pm 94$
TD-GALAXY (Zhang et al.,, 2022) $262\pm 5$ $1799\pm 125$
TD-Cluster Margin (Citovsky et al.,, 2021) $268\pm 9$ $6503\pm 82$
@@ D.3 Computational cost of TD-SPLIT and TD-FT
Table 6 shows the computational cost of running active learning for the TD-SPLIT and TD-FT approaches. We see that, overall, the TD-FT is significantly cheaper. This is a result of only being required to train on the unlabeled data whereas the TD-SPLIT approach requires training on both the unlabeled and labeled data simultaneously.
@@
Table 6: Total wall time in minutes of TD-SPLIT and TD-FT for THE F+MNSITS, CIFAR-10+100 and CheXpert datasets. We report the mean $\pm 1$ standard error over 4 seeds.
@@ D.4 Final test accuracies for Figures 3, 4
TD-SPLIT
EPIG $\mathbf{98.46\pm 0.17}$ $\mathbf{76.47\pm 0.27}$
@@
  TD-FT
EPIG $\mathbf{99.56\pm 0.10}$ $\mathbf{83.23\pm 0.38}$

### v2

@@ A.1.3 Models
For F+MNIST, we used the encoder and decoder from Burgess et al. (2017); for CheXpert we used the encoder and decoder from Higgins et al. (2017); for CIFAR-10+100 we used the ResNet–VAE used for the CIFAR-10 dataset in Kingma et al. (2016).
@@
For F+MNIST and CheXpert we used the ResNet18 architecture (He et al., 2016); for CIFAR-10+100 we used the ResNet50 architecture (He et al., 2016).
@@ A.1.4 Active learning
Initial labelled set:
To create our initial labelled set, we randomly sampled 2 labels from “0”, “1”, and “2” for F+MNIST; for CheXpert we randomly sampled 4 labels from “0” and “1”; for CIFAR-10+100 we randomly sampled 2 classes for each of our target classes and “0” class.
@@ A.1.5 Acquisition strategies
For our acquisition strategies, we used EPIG, BALD, Confidence Sampling and the acquisition strategies from our baselines methods which we describe below. To estimate BALD and EPIG, we used the same setup as Bickford Smith et al. (2023): we used 100 realisations of $\theta_{h}$ (for random forests this meant using the individual trees; otherwise this meant sampling from the parameter distribution). For EPIG, we used $M$ samples of $x_{*}$ from a finite set of unlabelled inputs representative of the downstream task, where $M=500$ for CheXpert and F+MNIST, and $M=1000$ for CIFAR-10+100.
@@
We used batch acquisitions for all datasets, with a batch size of 10 for F+MNIST and 100 for the others. We used the “power” batch acquisition strategy from (Kirsch et al., 2021) with $\beta=4$ for F+MNIST and $\beta=8$ otherwise. We make this choice as this strategy is both highly scalable and has been shown to give performance comparable to more sophisticated batch acquisition strategies.
@@ A.1.6 Baselines
We used baselines which were specifically designed for scenarios with class imbalance and redundancy (GALAXY (Zhang et al., 2022), SIMILAR (Kothawade et al., 2021)) or which have shown strong performance in such settings (Cluster Margin (Citovsky et al., 2021) from the results in Zhang et al. (2022)). We note that Confidence Sampling (Settles, 2009), though not treated as a baseline, has also shown strong performance in this setting (as per the results in Zhang et al. (2022)).
@@
For SIMILAR, we used the FLQMI relaxation of the submodular mutual information (SMI) due to memory constraints posed by the FLCMI relaxation. For Cluster Margin, we first experimented with their original hyperparameters ($\epsilon$ such that we have at least 10 clusters and $k_{m}=10k_{t}$). We found this to result in poor performance and so instead adopted the hyperparameters used in Zhang et al. (2022), i.e. choosing $\epsilon$ such that we have exactly 50 clusters and $k_{m}=1.25k_{t}$, where $k_{t}$ is our acquisition batch size.
@@
For a fair comparison with our approach, we used a ResNet18 model for all the baselines on F+MNIST and CheXpert, and a ResNet50 for CIFAR-10+100. We added a final fully–connected hidden layer of 128 hidden units on top of the ResNet models and trained them in a fully supervised fashion. For Cluster Margin, we followed Citovsky et al. (2021) and warm–started the models by training them on a validation set (0.5% of our pool size) that was evenly balanced between the target classes and “other” class; [^1] Note that Cluster Margin still performs worse than our approach despite the unrealistic assumption of a validation set. for SIMILAR, we followed Kothawade et al. (2021) and trained the models fully from scratch; for GALAXY, we followed Zhang et al. (2022) and initialised the ResNet backbones with weights pretrained on ImageNet in a fully supervised fashion. We trained the models using a batch size of 200, the Adam optimiser, and a learning rate of 0.001 for Cluster Margin, SIMILAR and 0.0001 for GALAXY.
@@
For the TRANSFER approach, we replaced our SimCLRv2 encoders pretrained on the pool with: ResNet50 and ResNet101 (He et al., 2016) pretrained on ImageNet in a fully supervised fashion, and DinoV2–Small and DinoV2–Big (Oquab et al., 2023) pretrained in a self–supervised fashion on the curated dataset discussed in Oquab et al. (2023) [^2] The ResNet models were accessed through TorchVision and the DinoV2 models were accessed from the PyTorch Hub (Paszke et al., 2019). Specifically, DinoV2–Small corresponds to facebookresearch/dinov2/dinov2_vits14 and DinoV2–Big corresponds to facebookresearch/dinov2/dinov2_vitb14..
@@
DinoV2 models:
Similar to (Oquab et al., 2023), we followed a similar finetuning setup to the one in Touvron et al. (2022). Specifically, we used a batch size of 512, the Adam optimiser, learning rate of $3\times 10^{-4}$ with 5 epochs warmup and cosine decay, and label smoothing with level 0.1. We used the same augmentations in Section A.1.2.
@@ Appendix B Additional Plots
B.1 Full Active Learning Curves for Table 1
@@
Figure 5: Test accuracy for our textbfTD-SPLIT approach on F+MNIST, CIFAR-10+100, CheXpert and the baselines considered in Table 1. All experiments were run for 4 seeds. Solid line shows mean and shading $\pm 1$ standard error.
@@
Figure 6: Test accuracy for our TD-FT approach on F+MNIST, CIFAR-10+100, CheXpert and the baselines considered in Table 1. All experiments were run for 4 seeds. Solid line shows mean and shading $\pm 1$ standard error.
@@ C.3 Different Prediction Heads
To show that our approach is compatible with different prediction heads, we replaced the random forest prediction head with a 1 layer neural network with 128 hidden units. We used the Laplace approximation (MacKay, 1992a) to infer the parameter distribution, where we used a standard Gaussian prior, $\mathcal{N}(0,I)$, and a diagonal, tempered posterior (Aitchison, 2020), with tempering implemented by raising the likelihood term to a power of $\text{dim}(\theta_{h})$ (i.e. the parameter count of the prediction head).
@@ Appendix D Additional Results
In this section, we provide additional results about the computational cost of our approach, using our task–driven representations for the baselines, and the acquisition counts for different approaches.
@@ D.1 Task–Driven Representations Improve Baselines
We see that integrating the baselines into our approach significantly boosts their performance on all datasets. We note, however, that their performance is still lagging behind our approach with the EPIG acquisition strategy.
@@
SIMILAR (Kothawade et al., 2021) $93.82\pm 0.18$ $30.87\pm 1.57$ $71.94\pm 0.47$
GALAXY (Zhang et al., 2022) $84.74\pm 0.74$ $55.28\pm 0.42$ $78.76\pm 0.35$
Cluster Margin (Citovsky et al., 2021) $94.24\pm 0.17$ $32.79\pm 0.45$ $75.41\pm 0.29$
US+EPIG (SimCLRv2, (Bickford Smith et al., 2024)) $98.53\pm 0.12$ $6.19\pm 0.42 77$$7.84\pm 0.28$
US+EPIG (VAE, (Bickford Smith et al., 2024)) $94.50\pm 0.34$ $30.47\pm 1.17$ $66.69\pm 0.56$
@@
TD-SIMILAR (Kothawade et al., 2021) $98.05\pm 1.19$ $9.90\pm 0.23 77$$7.26\pm 1.06$
TD-GALAXY (Zhang et al., 2022) $99.30\pm 0.13$ $73.62\pm 0.77$ $78.07\pm 0.73$
TD-Cluster Margin (Citovsky et al., 2021) $99.32\pm 0.18$ $70.47\pm 2.58$ $76.30\pm 0.38$
@@ D.2 Acquisition Counts for Different Approaches
From Table 6, we note two things. First we note that using our approach, whether that is with the baselines or instead of unsupervised representations, improves the count of the target classes by a large margin. This suggests that the gains displayed in Table 5 are not merely from using a better model, but also from making better acquisitions. Secondly, we note that the best performing method (TD-FT) does not have the largest amount of target classes acquired. This suggests that the precise data point acquired is important, not just its class (Yang et al., 2023).
@@
SIMILAR (Kothawade et al., 2021) $486\pm 6$ $920\pm 237$
GALAXY (Zhang et al., 2022) $118\pm 13$ $1806\pm 32$
Cluster Margin (Citovsky et al., 2021) $32\pm 1$ $1287\pm 53$
US+EPIG (SimCLRv2, (Bickford Smith et al., 2024)) $119\pm 8$ $936\pm 35$
US+EPIG (VAE, (Bickford Smith et al., 2024)) $118\pm 4$ $529\pm 17$
TD-SIMILAR (Kothawade et al., 2021) $493\pm 1$ $2231\pm 94$
TD-GALAXY (Zhang et al., 2022) $262\pm 5$ $1799\pm 125$
TD-Cluster Margin (Citovsky et al., 2021) $268\pm 9$ $6503\pm 82$
@@ D.3 Computational Cost of TD-SPLIT and TD-FT
Table 7 shows the computational cost of running active learning for the TD-SPLIT and TD-FT approaches. We see that, overall, TD-FT is significantly cheaper. This is a result of only being required to train on the labelled data whereas the TD-SPLIT approach requires training on both the unlabelled and labelled data simultaneously.
@@
Table 7: Total wall time in minutes of TD-SPLIT and TD-FT for the F+MNIST, CIFAR-10+100 and CheXpert datasets. We report the mean $\pm 1$ standard error over 4 seeds.
@@ D.4 Final Test Accuracies for Figures 3, 4
TD-SPLIT EPIG $\mathbf{98.46\pm 0.17}$ $\mathbf{76.47\pm 0.27}$
@@
  TD-FT EPIG $\mathbf{99.56\pm 0.10}$ $\mathbf{83.23\pm 0.38}$